# A Constant Factor Approximation for Directed Feedback Vertex Set in Graphs of Bounded Genus

### v3

@@ 1 Introduction
In the directed feedback vertex set problem (DFVS), we are given a directed graph $G=(V,E)$ and costs $c_{v}\ \forall v\in V$ and wish to find a minimum cost set $X$ for which $G\backslash X$ is acyclic. DFVS is one of Karp’s original 21 NP-hard problems [15]. The DFVS problem has many applications including deadlock resolution [10], VLSI chip design [18] and program verification [19].
A 2-approximation for (undirected) FVS is given in [2]. DFVS has a 2-approximation in tournaments [20] and bipartite tournaments [23], is polynomial-time solvable on graphs of bounded treewidth, has a 2.4-approximation in planar graphs [3] and has an $O(\log{n}\log{\log{n}})$-approximation in general graphs [7].
@@
$\min~{}~{}$ $c^{T}x$ ($\mbox{P}_{\text{DFVS}}$) s.t. $x(C)\geq 1\quad\forall\ C\in\mathcal{C}$ (1) $x\geq 0\enspace$ $\operatorname{max}~{}~{}$ $\mathbb{1}^{T}y$ ($\mbox{D}_{DFVS}$) s.t. $\sum_{C\in\mathcal{C},v\in C}y_{C}\leq c_{v}\quad\forall v\in V(G)$ $y\geq\mathbb{0}\enspace.$
@@
[…]
From the proof of Theorem 1 it is clear that the algorithm in Theorem 1 runs in time $O(g)\text{poly}(|V(G)|)$.
@@ 2 Preliminaries
Throughout this paper, all surfaces are orientable and smooth and all curves are piece-wise smooth. Distances on surfaces will refer to the geodesic (shortest path on the surface) distance. It is well known (see for instance [21]) that a smooth orientable surface $Q$ is diffeomorphic to the $g$-genus torus for some $g$. We henceforth assume our surface is a $g$-genus torus for some $g$.
Let us call a cycle of $G$, or a closed curve embedded on a surface facial if it bounds a region of the surface homeomorphic to the plane. It is well-known that a cycle is facial if and only if it partitions the surface into 2 connected components. For any set $F\subset V$, define $G^{F}$ to be the residual graph, that is, the subgraph of $G$ induced by those vertices that lie in a dicycle of $G\backslash F$.
3 Solving the case of no facial cycles
We first show the LP gap of the natural LP ($\mbox{P}_{\text{DFVS}}$) for $G$ has integrality gap $O(1)$ in the case $G$ contains no facial cycles. As mentioned previously facial cycles can be tackled in a similar way to that in [12].
Lemma 1.
[…]
@@ 3 Solving the case of no facial cycles
We prove the statement by induction on the genus $g$. The case $g=0$ was shown in [3]. Suppose the statement is true for $g=g^{\prime}$. Let $Q$ be a surface of genus $g$, Let $G$ be a digraph embedded on $Q$.
First, while the optimal solution $\bar{x}$ to ($\mbox{P}_{\text{DFVS}}$) has a vertex $v$ with value $\bar{x}_{v}\geq\frac{1}{24}$ add v to our temporary hitting set $F$. Formally initialize $F=\emptyset$. Obtain set $F\subset V$ as follows initially set $F=\emptyset$. While the optimal solution $\bar{x}$ to ($\mbox{P}_{\text{DFVS}}$) for $G^{F}$ contains a value $\bar{x}_{v}$ which is $\frac{1}{24}$ or more add $v$ to $F$.
We now seek to define distances on $G^{F}$. By standard LP theory, $\hat{x}$ has rational coordinates. Let $N\in\mathbb{Z}_{>0}$ be such that $N\hat{x}$ and $\frac{1}{12}N$ are integral, call $N\hat{x}_{v}$, the weight of $v$. For a subgraph $H$ of $G^{F}$ define the weighted distance $d_{\{}\omega,H\}(u,v)$ from $u$ to $v$ the minimum weight of a $u$-$v$ path in $H$. Define $d_{\omega}:=d_{\{\omega,G\}}$ Define $d_{\omega,H}(U,W):=\min_{u\in U,w\in W}\ d_{\omega}(u,w)$. Define $d_{\omega}(U,W)=d_{\omega,G}(U,W)$. Define the weighted distance of a closed walk $P^{\prime}=v_{0},v_{1},..,v_{l}v_{0}$, $\omega(P^{\prime})$ to be $\omega(P^{\prime}):=\sum_{i=0}^{l}N\hat{x}_{v_{i}}$. Since $\hat{x}$ is feasible the following result holds.
Proposition 2.
For any $W\subset V(H)$ they define “layers” $L_{i}=\{v\in H:\ \ d_{H}(W,v)=i\}$ the nodes at distance $i$ from $W$. They show that the cost of all layers $L_{0},L_{1},...$ is $\sum_{v\in V(G)}N\hat{x}_{v}$. This is very useful for us as we will use this to show that one layer in $L_{1},...L_{m}$ has cost at most $\frac{1}{m}\sum_{v\in V}N\hat{x}_{v}$. However the bypassing and replacing a node with a chain operation of [7] do not preserve the genus of the graph. We instead define the notion of weighted distance $d_{\omega}$. Denote the $i$-th layer from $W$ as $L_{i}:=\{v\in V:i\geq d_{\omega}(W,v)>i-\omega(v)\}$ the set of nodes for which the distance from $W$ to $v$ is at most $i$, but for which the distance plus the weight of $v$ is more than $i$. One can see that $v$ lies in $N\hat{x}_{v}$ different $L_{i}$, which is analogous to how $H$ defined in [7] contains $N\hat{x}_{v}$ copies of $v$ each lying in different layers as well. In particular a node of weight 0 does not lie in any $L_{i}$, which is analogous to how a vertex of weight 0 is bypassed in [7].
Consider the embedding of $C^{1}$ on our surface. Given a subgraph $W$ of $G$ denote by $g(W)$ the subset of our surface occupied by a vertex or edge of $W$. We want to define a “small” neighbourhood around $g(C^{1})$, not containing any vertices outside $C^{1}$, which we divide up into “left” of $g(C^{1})$ and “right” of $g(C^{1})$, which we do using the following propositions. These are slightly informal statements of the exact propositions we require, the precise statements appear in Appendix A.
Proposition 3.
Proposition 4.
We defer the proofs of 3 and 4 for now. We apply 3 to $g(C^{1})$ let $L,R$ be as in 3 so that each $g(e)$ for $e\in E(G)\backslash E(C^{1})$ is disjoint from at least one of $L,R$ and for each $e\in E(G\backslash C^{1})$, $g(e)$ is disjoint from both $L,R$.
Figure 1: $L$ and $R$ from 3 in yellow and red respectively curve $C^{\prime}$ depicted in black. The curve $h$ leaving $C^{\prime}$ from the left and entering from the right is depicted in dark green. The closed curve formed by $h$ and the subcurve of $C^{\prime}$ between $h(0)$ and $h(1)$ depicted in light green forms a non-facial closed curve.
Let $u^{\prime}_{i,1},u^{\prime}_{i,2},..,u^{\prime}_{i,l_{i}}$ be the out neighbours of $v_{i}$ such that the edges $u^{\prime}_{i,t^{\prime}}v_{i}$ reach $v_{i}$ from the left, that is, the arc obtained from reversing the arc $v_{i}u^{\prime}_{i,t^{\prime}}$ of our graph reaches $v_{i}$ from the left. Let $w^{\prime}_{i,1},w^{\prime}_{i,2},..,w^{\prime}_{i,z_{i}}$ the in neighbours of $v_{i}$ such that the edges $w^{\prime}_{i,t^{\prime}}v_{i}$ reach $v_{i}$ from the right. Subdivide each edge $v_{i}u^{\prime}_{i,t}$ into a path $v_{i}u_{i,t}u^{\prime}_{i,t}$ and each edge $w^{\prime}_{j,t^{\prime}}v_{j}$ into a path $w^{\prime}_{j,t^{\prime}}w_{j,t^{\prime}}v_{j}$ and give the new vertices $w_{j,t^{\prime}},u_{i,t}$ infinite cost. There is a natural embedding of our new graph on our surface by placing each $u_{i,t}$ where the midpoint of the curve $g(v_{i}u^{\prime}_{i,t})$ was embedded and likewise for $w_{j,t^{\prime}}$. By abuse of notation, we continue to call our graph $G$ and define $\hat{x}_{u_{i,t}}=\hat{x}_{w_{j,t^{\prime}}}=0$ for all $u_{i,t}$, $w_{j,t^{\prime}}$. Denote $U:=\cup_{i=1}^{l^{\prime}}\{u_{i,1},u_{i,2},..,u_{i,l_{i}}\}$ and $W:=\cup_{i=1}^{l^{\prime}}\{w_{i,1},w_{i,2},..,u_{i,l_{i}}\}$. For $X\subset[l^{\prime}]$, denote $U_{X}:=\cup_{i\in X}\{u_{i,1},u_{i,2},..,u_{i,l_{i}}\}$, $V_{X}=\{v_{i}:\ \ i\in X\}$ and $W_{X}:=\cup_{i\in X}\{w_{i,1},w_{i,2},..,w_{i,z_{i}}\}$.
Claim 1.
[…]
Proof.
Figure 2: Nodes of $U_{\tau_{+}}$ and $V_{\tau_{-}}$ shown in blue.
If $j\notin\tau_{-}$, then $w_{j,t^{\prime}}$ is not reachable from $U$ in $G^{F}\backslash S$. If $i\notin\tau_{+}$, then $W$ is not reachable from $u_{i,t}$ in $G^{F}\backslash S$. Thus if either $j\notin\tau_{-}$ or $i\notin\tau_{+}$ then there is no path from $u_{i,t}$ to $w_{j,t^{\prime}}$ in $G^{F}\backslash S$. Thus, $j\in\tau_{-}$ and $i\in\tau_{+}$. As $K^{\prime}$ is strongly connected, this implies that $G^{F}\backslash S$ contains a path from $V_{\tau_{-}}$ to $V_{\tau_{+}}$, which is not possible.
Now suppose that $d_{\omega}(V_{\tau_{-}},V_{\tau_{+}})\leq\frac{1}{12}N$. Let $i\in\tau_{-}$ and $j\in\tau_{+}$ be such that $d_{\omega}(v_{i},v_{j})\leq\frac{1}{12}N$ Let $P_{1},P_{2},P_{3}$ be $u_{a,t}$-$v_{i}$, $u_{j,t^{\prime}}$-$v_{b}$ and $v_{i}$-$v_{j}$ paths of weight at most $\frac{1}{12}N$, with the second last vertices of $P_{1},P_{2}$ being in $W$, for some $a,b$. Such paths exist as $i\in\tau_{-}$ and $j\in\tau_{+}$. If $a=i$, then $P_{1}v_{i}u_{a,t}$ is a cycle for which $\sum_{v\in P_{1}v_{i}u_{a,t}}\hat{x}_{v}<1$ which is a contradiction. So $a\neq i$, likewise $b\neq j$.
We claim $\omega(C^{1}_{(v_{a},v_{i})}),\omega(C^{1}_{(v_{j},v_{b})})\leq\frac{1}{12}N$. Suppose for a contradiction that $\omega(C^{1}_{(v_{a},v_{i})})>\frac{1}{12}N$ Since $\omega(C^{1})=N$, this implies that $\omega(C^{1}_{(v_{i+1},v_{a-1})})<N-\frac{1}{12}N$. Then the cycle $P_{1}C^{1}_{(v_{i},v_{a})}v_{a}u_{a,t}$ satisfies $\sum_{v\in V(P_{1}C^{1}_{(v_{i},v_{a})}v_{a}u_{a,t})}\hat{x}_{v}<1-\frac{1}{12}+\frac{1}{12}=1$ which is a contradiction. Likewise $\omega(v_{j},v_{b})\leq\frac{1}{12}N$.
Now suppose that $i=j$. Suppose for a contradiction that $C^{1}_{(v_{a},v_{i})}\cap C^{1}_{(v_{i},v_{b})}\neq\{v_{i}\}$. Let $v\in(C^{1}_{(v_{a},v_{i})}\cap C^{1}_{(v_{i},v_{b})})\backslash v_{i}$. Let $Q_{1}$ be a path from $v$ to $v_{i}$ in $C^{1}_{(v_{a},v_{i})}$ and $Q_{2}$ a path from $v_{i}$ to $v$ in $C^{1}_{(v_{i},v_{b})}$. Then $Q_{1}Q_{2}$ is a closed walk of weighted distance at most $\frac{1}{6}N$, which is a contradiction.
Claim 2.
Proof.
Claim 3.
Proof.
Recall that any dicycle of $G^{F}\backslash R^{\prime}$ is contained in either $G^{F}\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ or $G^{F}\backslash(P_{2}\cup C^{1}_{(v_{j+1},v_{b})})$. Since $v_{j}$ is not contained in any dicycle of $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$ it follows that any dicycle of $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$ is either contained in $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ or in $G\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$. By 4, $g(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ and $P_{2}\cup C^{1}_{(v_{j},v_{b})}$ are nonfacial.
Definition 1.
[1, 13] Given a simple closed curve $f$ on a surface without boundary $Q$, not dividing the surface into 2 regions, we say $Q^{\prime}$ is obtained by doing surgery along $f$ if $Q^{\prime}$ is obtained as follows. “Thicken” $f$ to obtain a cylinder and remove this cylinder from $Q$, call this resulting surface $Q^{\prime\prime}$. The boundary of $Q^{\prime\prime}$ consists of 2 circles we “glue” two cones along these circles and call this final surface $Q^{\prime}$.
Theorem 5.
([1] p.162) For a surface without boundary $Q$ of genus $g^{\prime}$, $Q^{\prime}$ obtained by 1 is a surface without boundary of genus at most $g^{\prime}-1$.
We apply the surgery of 1 to $g(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ to obtain a surface $Q^{\prime}$ of genus one less than $Q$. Thus, $G^{F}\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ is of genus at least one less than $G^{F}$. Likewise $G^{F}\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$ is of genus at least one less than $G^{F}$. Now let $G_{1},G_{2},..,G_{l}$ be the strongly connected components of $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$. Since any closed walk of of $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$ is either contained in $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ or in $G\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$ each strongly connected component is contained in either contained in $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ or in $G\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$. Since both $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ and $G\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$ are of smaller genus, each $G_{i}$ is of smaller genus. By induction, there are solutions $A_{i}$ to $G_{i}$ of cost $c_{g-1}OPT_{LP(G_{i})}$, where $c_{g}$ is the integrality gap of the DFVS LP for graphs of genus $g$.
Define $\hat{x}^{G}_{i}\in\mathbb{R}^{V(G_{i})}$ as $\hat{x}^{G}_{i}(v)=\hat{x}_{v}$, where $\hat{x}$ is as in the proof of 1. Define $\hat{x}^{G}_{i}\in\mathbb{R}^{V(G_{i})}$ as $\hat{x}^{G}_{i}(v)=\hat{x}_{v}$, where $\hat{x}$ is as in the proof of 1. Since graphs $G_{i}$ are vertex disjoint, $\sum_{i=1}^{l}OPT_{LP(G_{i})}\leq\sum_{i=1}^{l}\sum_{v\in V(G_{i})}\hat{x}^{G}_{i}(v)\leq\sum_{v\in V(G)}\hat{x}=OPT_{LP(G)}$. Now $F\cup R^{\prime}\cup K^{\prime+}\cup(\cup_{i=1}^{l}A_{i})$ is a DFVS of cost $(O(1)+c_{g-1})OPT_{LP(G)}=(O(1)+O(g-1))OPT_{LP(G)}=(O(g))OPT_{LP(G)}$.
Let $b^{\prime}_{i,1},b^{\prime}_{i,2},..,b^{\prime}_{i,l^{\prime}_{i}}$ be the in neighbours of $v_{i}$ such that each edge $b^{\prime}_{i,t}v_{i}$ reaches $v_{i}$ from the left and $d^{\prime}_{i,1},d^{\prime}_{i,2},..,d^{\prime}_{i,t^{\prime}_{i}}$ be the out neighbours of $v_{i}$ such that the edge $d^{\prime}_{i,t^{\prime}}v_{i}$ reaches $v_{i}$ from the right. Subdivide each edge $b^{\prime}_{j,t^{\prime}}v_{j}$ into a path $b^{\prime}_{j,t^{\prime}}b_{j,t^{\prime}}v_{j}$ and each edge $v_{i}d^{\prime}_{i,t}$ into a path $v_{i}d_{i,t}d^{\prime}_{i,t}$ and give the new vertices $d_{j,t^{\prime}},b_{i,t}$ infinite cost. There is a natural embedding of our new graph on our surface by placing each $b_{i,t}$ where the midpoint of the curve $g(v_{i}b^{\prime}_{j,t^{\prime}})$ was embedded and likewise for $d_{j,t^{\prime}}$. By abuse of notation, we continue to call our graph $G$ and define $\hat{x}_{b_{i,t}}=\hat{x}_{d_{j,t^{\prime}}}=0$ for all $b_{i,t}$, $d_{j,t^{\prime}}$.
Claim 4.
[…]
Proposition 6.
If there is a (undirected) path $P=q_{1},q_{2},..,q_{t}w_{i,t}$ from some $w_{i,t}\in W$ (resp $b_{i,t}\in B$) in $\operatorname{un}(K)\backslash C^{1}$, then there is a directed path from $U$ (resp to $q_{j}$ in $G^{F}\backslash(S\cup T\cup C^{1})$ for any $j=1,2,..,t$.
Proof.
Let $P=u_{i,t}q_{1},q_{2},..,q_{t}$ be a path in $\operatorname{un}(K)\backslash C^{1}$ from some $u_{i,t}\in U$ (resp $u_{i,t}\in D$). We prove by induction $t^{\prime}$ on that there is a directed path from $U$ to $q_{j}$ in $G^{F}\backslash(S\cup T\cup C^{1})$ for any $j=1,2,..,t^{\prime}$. The case $t^{\prime}=1$ is clear as each $u_{i,t}\in U$ (resp $u_{i,t}\in D$) only has out neighbours so the undirected edge $\{u_{i,t},q_{1}\}$ in $\operatorname{un}(K)$ is directed from $u_{i,t}$ to $q_{1}$.
Now assume the statement true for $t^{\prime}=t^{\prime\prime}$. For $t^{\prime}=t^{\prime\prime}+1$, if the undirected edge $\{q_{t^{\prime}},q_{t^{\prime}+1}\}$ is directed from $q_{t^{\prime}}$ to $q_{t^{\prime}+1}$, then there is a directed path from $u_{i,t}$ to $q_{j}$ in $G^{F}\backslash(S\cup T\cup C^{1})$.
Otherwise $\{q_{t^{\prime}},q_{t^{\prime}+1}\}$ is directed from $q_{t^{\prime}+1}$ to $q_{t^{\prime}}$. By strong connectedness of $K$, there is a directed path $P^{\prime}$ from $q_{t^{\prime}}$ to $q_{t^{\prime}+1}$ in $K\backslash(S\cup T)$. $P^{\prime}$ must intersect $C^{1}$. Thus $P^{\prime}$ must intersect $W$ or $B$. Let $P^{\prime\prime}$ denote the subpath of $P^{\prime}$ from $u_{i,t}$ to when $P^{\prime}$ first intersects $U$ or $B$. By construction $P^{\prime\prime}$ lies in $G^{F}\backslash(S\cup T\cup C^{1})$. As $u_{i,t}$ lies in $U$ (resp $D$) $P^{\prime\prime}$ does not intersect $W$ (resp. $B$). Thus $P^{\prime\prime}$ is a $u_{i,t}$-$B$ (resp. $u_{i,t}$-$W$) path. Consider the subpath $Q$ of the reversal of $P^{\prime}$ starting from $q_{t^{\prime}+1}$ to when the reversal of $P^{\prime}$ first intersects $D$ or $U$. Let $\operatorname{rev}(Q)$ denote the reversal of $Q$. Note $\operatorname{rev}(Q)$ lies in $G^{F}\backslash(S\cup T\cup C^{1})$. If the starting vertex of $\operatorname{rev}(Q)$ is in $D$ (resp. $U$), then $\operatorname{rev}(Q)\cup P^{\prime}$ is a $D$-$B$ (resp. $U$-$W$) path in $G^{F}\backslash(S\cup T\cup C^{1})$. This contradicts 4. Thus the starting vertex of of $\operatorname{rev}(Q)$ is in $U$ (resp. $D$). This implies there is a path from $U$ (resp. $D$ ) to $q_{t^{\prime}}$ completing the induction. The proof of the second part is similar. ∎
Proposition 7.
Proof.
If we have a $U$-$W$ path $P=u_{i,t}q_{1},q_{2},..,q_{t}w_{j,t^{\prime}}$ in $\operatorname{un}(K)\backslash C^{1}$, then by 6, there are directed $U$-$q_{1}$ and $q_{1}$-$W$ paths $P_{1}$ and $P_{2}$ in $\operatorname{un}(K)\backslash C^{1}$. Then $P_{1}\cup P_{2}$ is a directed $U$-$W$ path in $K\backslash C^{1}$ which contradicts 4. Thus we do not have a $U$-$W$ path $P=u_{i,t}q_{1},q_{2},..,q_{t}w_{j,t^{\prime}}$ in $K\backslash C^{1}$. Likewise we do not have a $D$-$B$ path $P=u_{i,t}q_{1},q_{2},..,q_{t}w_{j,t^{\prime}}$ in $K\backslash C^{1}$.
Suppose we have a $U$-$D$ path $P=u_{i,t}q_{1},q_{2},..,q_{t}w_{j,t^{\prime}}$ in $\operatorname{un}(K)\backslash C^{1}$. By 6, there are directed $U$-$q_{1}$ and $D$-$q_{1}$ paths $P_{1}$ and $P_{2}$ in $K\backslash C^{1}$. Recall $U$ has no in-neighbours of in $G\backslash C^{1}$, so the edge $\{u_{i,t},q_{1}\}$ in $K$ is directed from $u_{i,t}$ to $q_{1}$. By 2 connectedness of $K$, there is a path $P_{3}$ from $q_{1}$ to $u_{i,t}$. The only in-neighbours of $u_{i,t}$ are in $C^{1}$, thus $P_{3}$ intersects $W\cup B$. Let $P^{\prime}_{3}$ be the subpath of $P_{3}$ from $q_{1}$ to when it the path first intersects $W\cup B$. If the endpoint of $P^{\prime}_{3}$ is in $W$, then $P_{1}\cup P^{\prime}_{3}$ is a $U$-$W$ path in $K\backslash C^{1}$. Otherwise, if the endpoint of $P^{\prime}_{3}$ is in $B$, then $P_{2}\cup P^{\prime}_{3}$ is a $D$-$B$ path in $K\backslash C^{1}$. Either way this contradicts 4.
Proposition 8.
[14][24]Suppose $G$ is a graph embedded on a surface $Q$. Let $C$ be a cycle of $G$ that does not divide $Q$ into two separate regions such that there is no edge between vertices of $C$ that is not part of $C$. Define a “left and right as in [] Let $\hat{L}$, $\hat{R}$ denote the neighbours of $C$ that are “left” or “right” of $C$. Suppose each connected component of $G\backslash C$ only contains nodes of $\hat{L}$ or $\hat{R}$ but not both. Then $G$ is embeddable on a surface of smaller genus.
Applying 7, we get that each strongly connected component $K$ of $G^{F}\backslash C^{1}$ satisfies 8. Thus $K$ is embeddable on a surface of $g-1$. Let $G_{1},G_{2},.,G_{l}$ be the strongly connected components of $G^{F}\backslash(S\cup T)$, so each $G_{i}$ is embeddable on a surface of $g-1$. By induction the LP gap for any strongly connected component $K$ of $G^{F}\backslash C^{1}$ is at most some constant $c_{g-1}$. By induction, there are solutions $A_{i}$ to $G_{i}$ of cost $c_{g-1}OPT_{LP(G_{i})}$, where $c_{g}$ is the integrality gap of the DFVS LP for graphs of genus $g$. Define $\hat{x}^{G}_{i}\in\mathbb{R}^{V(G_{i})}$ as $\hat{x}^{G}_{i}(v)=\hat{x}_{v}$, where $\hat{x}$ is as in the proof of 1. Since graphs $G_{i}$ are vertex disjoint, $\sum_{i=1}^{l}OPT_{LP(G_{i})}\leq\sum_{i=1}^{l}\sum_{v\in V(G_{i})}\hat{x}^{G}_{i}(v)\leq\sum_{v\in V(G)}\hat{x}=OPT_{LP(G)}$. Then $S\cup T\cup F\cup(\cup_{i=1}^{l}A_{i})$ is a DFVS of cost $(O(1)+c_{g-1})OPT_{LP(G)}=(O(1)+O(g-1))OPT_{LP(G)}=(O(g))OPT_{LP(G)}$.
As observed in [7], ($\mbox{P}_{\text{DFVS}}$) can be solved in polynomial-time via the ellipsoid method. Hence 1 yields a polynomial time $O(g)$-approximation algorithm for DFVS in graphs of genus $g$ with no facial cycle.
4 Solving the general case
We now show how to obtain an $O(g)$-approximation for DFVS for graphs embedded on any surface $Q$ of genus $g$.
See 1
Proof.
@@ 4 Solving the general case
We prove the statement by induction on the genus $g$. The case $g=0$ was shown in [3]. Suppose the statement is true for $g=g^{\prime}$. Let $Q$ be a surface of genus $g$. Let $G$ be a digraph embedded on $Q$. If $C$ is a facial cycle bounding a face of $G$ call $C$ face-minimal. Note that if $G$ contains a facial cycle, then it must contain a face minimal cycle by the following argument. We associate each facial cycle $C$ with the region $R$ of $Q$ homeomorphic to the plane that it bounds. Let $C$ be a dicycle bounding a minimal (by containment) region of $Q$. If $C$ is not a face then there is a dipath $P$ between two nodes $u,v$ of $C$ then $P$ together with either the $u$-$v$ or $v$-$u$ dipath in $C$ forms a cycle bounding a smaller region of $Q$, which is a contradiction.
@@
Our algorithm proceeds as follows. The first part is a primal-dual algorithm analogous to the technique of [12] for DFVS in planar graphs. Given feasible dual solution $y$ to ($\mbox{D}_{DFVS}$), let the residual cost of node $v\in V$ be $c_{v}-\sum_{C\in\mathcal{C},v\in C}y_{C}$. For $\hat{S}\subset V(G)$, recall $G^{\hat{S}}$ denotes the subgraph of $G$ induced by those vertices which are in a dicycle of $G\backslash\hat{S}$.
@@
In the second step using 1, we find a DFVS $\tilde{S}$ of $G^{\bar{S}}$ of cost at most $O(g)OPT_{LP}$.

### v4

@@ 1 Introduction
In the directed feedback vertex set problem (DFVS), we are given a (node-weighted) directed graph $G=(V,E)$ with costs $c_{v}\ \forall v\in V$ and wish to find a minimum cost set $X$ for which $G\backslash X$ is acyclic. DFVS is one of Karp’s original 21 NP-hard problems [16]. The DFVS problem has many applications including deadlock resolution [10], VLSI chip design [19] and program verification [20].
A 2-approximation for (undirected) FVS is given in [2]. DFVS has a 2-approximation in tournaments [21] and bipartite tournaments [24], is polynomial-time solvable on graphs of bounded treewidth, has a 2.4-approximation in planar graphs [3] and has an $O(\log{n}\log{\log{n}})$-approximation in general graphs [7]. DFVS does not have an $O(1)$-approximation under the unique games conjecture [13]. The genus of a graph is the minimal integer $g$ such that the graph can be drawn without crossing itself on a sphere with $g$ handles.
@@
$\min~{}~{}$ $c^{T}x$ ($\mbox{P}_{\text{DFVS}}$)
s.t. $x(C)\geq 1\quad\forall\ C\in\mathcal{C}$ (1)
$x\geq 0\enspace$
$\operatorname{max}~{}~{}$ $\mathbb{1}^{T}y$ ($\mbox{D}_{DFVS}$)
s.t. $\sum_{C\in\mathcal{C},v\in C}y_{C}\leq c_{v}\quad\forall v\in V(G)$
$y\geq\mathbb{0}\enspace.$
@@
For any fixed genus $g$, there is a polynomial-time $O(g)$-approximation for DFVS for graphs of genus $g$. Moreover, the algorithm returns a DFVS with cost $O(g)$ times the optimum solution to ($\mbox{P}_{\text{DFVS}}$).
From the proof of Theorem 1, it is clear that the algorithm in Theorem 1 runs in time $O(g)\text{poly}(|V(G)|)$.
@@ 2 Preliminaries
Throughout this paper, all surfaces are orientable and smooth and all curves are piece-wise smooth. Distances on surfaces will refer to the geodesic (shortest path on the surface) distance. It is well known (see for instance [22]) that a smooth orientable surface $Q$ is diffeomorphic to the $g$-genus torus for some $g$. For $X\subset Q$, denote $\operatorname{cl}(X)$ as the closure of $X$ in $Q$. We henceforth assume our surface is a $g$-genus torus for some $g$.
Let us call a cycle of $G$, or a closed curve $C$ embedded on a surface $Q$ facial if it bounds a region $\operatorname{inside}(C)$ of the surface homeomorphic to the plane. If the genus of $Q$ is greater than 0, call $\operatorname{inside}(C)$ the inside region of $Q\backslash C$. For any set $F\subset V$, define $G^{F}$ to be the residual graph, that is, the subgraph of $G$ induced by those vertices that lie in a dicycle of $G\backslash F$.
3 Hitting the facial cycles of a digraph
In general, given a (node-weighted) directed graph $G=(V,E)$ with costs $c_{v}\ \forall v\in V$ a set $\mathcal{C}$ of cycles of a digraph $G$, we define the $\mathcal{C}$-hitting set problem as the problem of finding a minimum cost set $X$ such that $X\cap C\neq\emptyset\ \forall C\in\mathcal{C}$. In this section, we are concerned with when $\mathcal{C}$ is the set of facial cycles of our graph.
Given a digraph $G$ embedded on a surface $Q$, We show how to obtain an $O(g)$-approximation for the problem of finding a minimal hitting set for the set of facial dicycles of a graph embedded on any surface $Q$ of genus $g$.
Theorem 2.
For a graph $G$ embedded on a surface $Q$ of genus $g$, there is a polynomial-time $O(g)$-approximation for the problem of hitting facial dicycles of $G$. Moreover, the algorithm returns a DFVS with cost $O(g)$ times the optimum solution to ($\mbox{P}_{\text{DFVS}}$) where $\mathcal{C}$ is the set of facial dicycles.
Further, if $G$ is embedded in a way such that there exist regions $R_{1},R_{2}$ of $Q$ homeomorphic to the open disk, such that the inside region of any facial dicycle contains at least one of $R_{1},R_{2}$, then there is an algorithm that returns the optimum solution to ($\mbox{P}_{\text{DFVS}}$) where $\mathcal{C}$ is the set of facial dicycles.
@@ 3 Hitting the facial cycles of a digraph
We first show that DFVS has an $O(g)$-approximation. If $C$ is a facial cycle bounding a face of $G$, call $C$ face minimal. Note that if $G$ contains a facial cycle, then it must contain a face minimal cycle by the following argument. Let $C$ be a dicycle such that $\operatorname{inside}(C)$ is a minimal (by containment) region of $Q$. Recall that we removed all vertices of $G$ not lying on a dicycle. In particular, any vertex $w$ inside the region $\operatorname{inside}(C)$ must lie on a facial dicycle $A_{w}$. $A_{w}$ cannot be contained entirely in $\operatorname{cl}(\operatorname{inside}(C))$, as then the region $A_{w}$ would be strictly contained in $\operatorname{inside}(C)$. Thus, $\operatorname{inside}(A_{w})$ intersects $C$ and there is a dipath $P$ between two nodes $u,v$ of $C$. If $C$ is not a face, then either there is a vertex $w$ inside the region $R_{C}$ or there is an edge $uv$ between two nodes $u,v$ of $C$ such that $g(uv)\backslash(g(v)\cup g(u))$ lies in $R_{C}$. In either case, there is a dipath $P$ between two nodes $u,v$ of $C$ then $P$ together with either the $u$-$v$ or $v$-$u$ dipath in $C$ forms a cycle bounding a smaller region of $Q$, which is a contradiction.
@@
Our algorithm proceeds as follows. This is a primal-dual algorithm analogous to the technique of [12] for DFVS in planar graphs. Given a feasible dual solution $y$ to ($\mbox{D}_{DFVS}$), let the residual cost of node $v\in V$ be $c_{v}-\sum_{C\in\mathcal{C},v\in C}y_{C}$. For $\hat{S}\subset V(G)$, recall $G^{\hat{S}}$ denotes the subgraph of $G$ induced by those vertices which are in a dicycle of $G\backslash\hat{S}$.